OpenStudy (anonymous):

help please Find the height of a square pyramid that has a volume of 75 cubic feet and a base length of 5 feet. (1 points) 5 feet 9 feet 25 feet 45 feet

OpenStudy (tkhunny):

You will need a formula for the Volume of a square pyramid.

OpenStudy (anonymous):

or nah I aint bout that life

OpenStudy (tkhunny):

Try to pay attention. You will need a formula for the Volume of a square pyramid.

OpenStudy (anonymous):

try actually helping

OpenStudy (tkhunny):

I am helping. You're trying not to learn. Time to start trying! All about pyramids. http://www.mathsisfun.com/geometry/pyramids.html Go find that formula!

OpenStudy (anonymous):

The volume=base (sq) x h/3 so use that formula to help

OpenStudy (tkhunny):

Wow!! The formula just dropped out of the sky. Okay, now how do you use it?

OpenStudy (anonymous):

i got the answer 25

OpenStudy (anonymous):

what is the base, first of all?

OpenStudy (anonymous):

base is 75

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the volume is 75

OpenStudy (anonymous):

fomula sayss volume = base and the volume is 75 so the ase must be 75

OpenStudy (anonymous):

base is 5

OpenStudy (anonymous):

good

OpenStudy (anonymous):

so what is 5(sq) or 5x5?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

yep. so our variable is the lenght, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this mean that the base,5, times the height, divided by three should equal 75

OpenStudy (anonymous):

75=5 x h/3

OpenStudy (anonymous):

so 45 is the correct answer is 45

OpenStudy (anonymous):

no

OpenStudy (anonymous):

45 x 5 =225/3=75

OpenStudy (anonymous):

start with 1, and if that's not the answer, go to 2, and so on.

OpenStudy (anonymous):

the answer is 45 bud i just did the math.

OpenStudy (anonymous):

either 45 or 233 is the height, it can't be both of them

OpenStudy (anonymous):

u said base (5) times height (45) divided by 3 should = 75 and it does so 45 is the correct answer

OpenStudy (anonymous):

So like I said start with 1 and keep going if 1 or 2 or 3 isn't the height

OpenStudy (anonymous):

this is multiple choice u know tht right? i dont have to do that.

OpenStudy (anonymous):

i used commen sense and picked the highest number used your formula and got the height which is 45

OpenStudy (anonymous):

no. start with the lowest number, 1

OpenStudy (anonymous):

nah im coorect i got a 100%

OpenStudy (anonymous):

I know 45 makes sense but that's not it.

OpenStudy (anonymous):

oh wow. good job

OpenStudy (anonymous):

thank you for the help

OpenStudy (anonymous):

nah. I didn't help ( I only got the formula)

OpenStudy (anonymous):

but thanks for the medal anyway